Ranga Ramasesh is the operations manager for a firm that is trying to decide which one of four countries it should research for possible outsourcing providers. The first step is to select a country based on cultural risk​ factors, which are critical to eventual business success with the provider. Ranga has reviewed outsourcing provider directories and found that the four countries in the table below have an ample number of providers from which they can choose. To aid in the country selection​ step, he has enlisted the aid of a cultural​ expert, John​ Wang, who has provided ratings of the various criteria in the table below. The resulting ratings are on a 1 to 10​ scale, where 1 is a low risk and 10 is a high risk.​ Further, John Wang has determined six criteria weightings. ​ Trust, with a weight of 0.40​; Quality, with 0.20​; ​Religious, with 0.10​; Individualism, with 0.10​; ​Time, with 0.10​; and​ Uncertainty, with 0.10. Using the weighted​ factor-rating method, which country should Ranga​ select?

In the following​ table, compute the total weighted score for each candidate ​(enter each response as a decimal rounded to one​ place).

                                                                                                                                  

Selection Criterion

Weight

Mexico

Panama

Costa Rica

Peru

Trust

0.40

4

1

3

2

Society value of quality work

0.20

6

4

8

8

Religious attitudes

0.10

3

3

3

5

Individualism attitudes

0.10

5

2

4

8

Time orientation attitudes

0.10

4

6

7

3

Uncertainity avoidance attitude

0.10

3

2

4

2

Total Weighted​ Score:

?

?

?

?
Answer #1

Weighted Score of Mexico = (0.4*4 + 0.2*6 + 0.1*3 + 0.1*5 + 0.1*4 + 0.1*3)

= (1.6 + 1.2 + 0.3 + 0.5 + 0.4 + 0.3)

= 4.3

Weighted Score of Panama = (0.4*1 + 0.2*4 + 0.1*3 + 0.1*2 + 0.1*6 + 0.1*2)

= (0.4 + 0.8 + 0.3 + 0.2 + 0.6 + 0.2)

= 2.5

Weighted Score of Costa Rica = (0.4*3 + 0.2*8 + 0.1*3 + 0.1*4 + 0.1*7 + 0.1*4)

= (1.2 + 1.6 + 0.3 + 0.4 + 0.7 + 0.4)

= 4.6

Weighted Score of Peru = (0.4*2 + 0.2*8 + 0.1*5 + 0.1*8 + 0.1*3 + 0.1*2)

= (0.8 + 1.6 + 0.5 + 0.8 + 0.3 + 0.2)

= 4.2

Based on the weighted score, Panama should be selected based on the lowest score which would mean lowest risk accordingly.